OpenStudy (anonymous):

Write the ratio in simplest form. 50:10 a) 25:5 b) 10:2 c) 5:1 d) 1:5 Solve the following proportions. (i'll put the equation in the chat thing)

OpenStudy (anonymous):

If we divided 50 on 10 what will we have ?;)

OpenStudy (anonymous):

Obviously 5:1

OpenStudy (charlotte123):

divide it by 10 on both sides and ya get that the answer is 5:1 :)

OpenStudy (anonymous):

(equation for solve the following proportions) \[\frac{ p }{ 6 }\]=\[\frac{ 7 }{ 8 }\]

OpenStudy (charlotte123):

p = 21/4 (5.25)